# MnLargeSymbols’164 MnLargeSymbols’171

### v2

@@ A.1 Adaptive Sampling Methods
Adaptive sampling methods have also been extended to tackle problems outside the unconstrained optimization setting. Xie et al. (2023) proposed a proximal extension of the norm test and the inner product test to minimize a convex composite objective, which comprises a stochastic function and a deterministic, potentially nonsmooth function. Beiser et al. (2023) explored deterministic constrained problems, including those with nonconvex objectives. Moreover, adaptive sampling methods have been applied to various optimization problems and algorithms, such as (Bollapragada et al., 2018b) for L-BFGS, (Berahas et al., 2022) for sequential quadratic programming (SQP) in the context of equality-constrained problems, (Bollapragada et al., 2023) for augmented Lagrangian methods, and (Bollapragada and Wild, 2023) for quasi-Newton methods.
@@ A.2 Convergence of Stochastic Gradient Methods
We give a brief overview on the convergence results of SGD and adaptive gradient methods. The convergence of SGD for smooth nonconvex functions was first analyzed in (Ghadimi and Lan, 2013), assuming a uniform bound on the variance of stochastic gradients. Arjevani et al. (2023) gave a tight lower bound. Bottou et al. (2018) extended the convergence result with the so-called affine variance noise model. Different assumptions on the second moment of stochastic gradients appearing in the literature were reviewed in (Khaled and Richtárik, 2023), with a view to developing a general convergence theory of SGD for smooth nonconvex functions.
@@ Appendix B The Overall Algorithm
Sample an i.i.d. data batch $\mathcal{B}_{k}\subseteq\mathcal{B}$ uniformly from $\mathcal{B}$ of size $b_{k}$
@@
$\mathsf{T}_{j}(x_{k};\mathcal{B}_{k},\eta)\coloneqq\frac{1}{b_{k}-1}\sum_{i\in\mathcal{B}_{k}}\left(\partial_{j}f(x_{k};\xi_{i})-\partial_{j}F_{\mathcal{B}_{k}}(x_{k})\right)^{2}\big{/}\left(\eta^{2}(\partial_{j}F_{\mathcal{B}_{k}}(x_{k}))^{2}\right),\;j\in\llbracket d\rrbracket$
@@ C.2 Preparatory Definitions, Propositions and Lemmas
This is part of the result of (Bollapragada et al., 2018a), Lemma 3.1. We include its proof here with our notation for completeness.
@@ C.2.3 Technical Lemmas
We state without proof a similar descent lemma for $(L_{0},L_{1})$-smooth functions. Its proof can be found in (Zhang et al., 2020a), Lemma A.3.
@@ C.3 Convergence Results for AdaGrad-Norm and AdaGrad
$\mathbb{E}_{k}[F(x_{k+1})]\\ \leqslant F(x_{k})-\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)g_{k}\right]\right\rangle+\frac{L\alpha^{2}}{2}\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}.$ (C.5)
@@
$\frac{L\alpha^{2}}{2}\sum_{k=1}^{K}\mathbb{E}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}+\frac{\tau L^{2}\alpha^{3}}{2}(1+\rho\tau)\sum_{k=1}^{K}\mathbb{E}\left[\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}\right]\\ \leqslant\frac{L\alpha^{2}}{2}(\mathbb{E}\log v_{K}-\log v_{0})+\tau L^{2}\alpha^{3}(1+\rho\tau)\frac{1}{\sqrt{v_{0}}}.$
@@
$\frac{\alpha}{2}\left(1-\frac{1}{\rho}\right)\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]\leqslant F(x_{1})-\mathbb{E}[F(x_{K+1})]+\frac{\tau\alpha}{2}\mathbb{E}[\varphi_{0}-\varphi_{K}]\\ +\frac{L\alpha^{2}}{2}(\mathbb{E}\log v_{K}-\log v_{0})+\tau L^{2}\alpha^{3}(1+\rho\tau)\frac{1}{\sqrt{v_{0}}},$
@@
By Lemma C.2, we have $\mathbb{E}[\sqrt{v_{k}}]\leqslant 2c_{1}\tau+8c_{2}\tau\log(c_{2}\tau+1)\eqqcolon c_{3}$.
Since, by Hölder’s inequality,
$c_{1}+2c_{2}\log\mathbb{E}[\sqrt{v_{k}}]\geqslant\sum_{k=1}^{K}\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]\\ \geqslant\mathbb{E}\left[\frac{1}{\sqrt{v_{K}}}\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}\right]\geqslant\frac{\mathbb{E}\left[\sqrt{\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}}\right]^{2}}{\mathbb{E}[\sqrt{v_{K}}]},$
@@
$\mathbb{E}\left[\sqrt{\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}}\right]^{\negthickspace 2}\leqslant\mathbb{E}[\sqrt{v_{K}}]\cdot\left(c_{1}+2c_{2}\log\mathbb{E}[\sqrt{v_{K}}]\right)\leqslant c_{3}(c_{1}+2c_{2}\log c_{3}).$
@@
As will be seen later, under the choice of the parameters $(\rho_{1},\rho_{2},\omega)$, the constant step size satisfies $\alpha\leqslant 1/L_{1}$ and $\left\lVert\frac{\nabla F(x_{k})}{\sqrt{v_{k}}}\right\rVert\leqslant 1$, we have
$\left\lVert x_{k+1}-x_{k}\right\rVert=\alpha\left\lVert\frac{\nabla F(x_{k})}{\sqrt{v_{k}}}\right\rVert\leqslant 1/L_{1}.$
@@
$\quad\;-\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle+\frac{L_{1}\alpha^{2}\left\lVert\nabla F(x_{k})\right\rVert}{2}\cdot\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}$ (C.13)
@@
since $\alpha\leqslant 1/L_{1}$.
By the A.M.-G.M. inequality (i.e., $2ab\leqslant\frac{1}{\rho_{1}}a^{2}+\rho_{1}b^{2}$ for any $\rho_{1}>0$), we have
@@
$\rho_{1}\alpha\tau\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right]\leqslant\alpha\left(\frac{\rho_{1}}{\rho_{2}}+2\omega\right)\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{2L_{0}^{2}}{\sqrt{v_{k-1}}}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}\\ +\frac{\rho_{1}\rho_{2}L_{0}^{2}\tau^{2}}{\sqrt{v_{k-1}}}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+\rho_{1}\alpha\tau\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}].$

### v3

@@ A.1 Adaptive Sampling Methods
Adaptive sampling methods have also been extended to tackle problems outside the unconstrained optimization setting. Xie et al. (2023) proposed a proximal extension of the norm test and the inner product test to minimize a convex composite objective, which comprises a stochastic function and a deterministic, potentially nonsmooth function. Beiser et al. (2023) explored deterministic constrained problems, including those with nonconvex objectives. Moreover, adaptive sampling methods have been applied to various optimization problems and algorithms, such as Bollapragada et al. (2018b) for L-BFGS, Berahas et al. (2022) for sequential quadratic programming (SQP) in the context of equality-constrained problems, Bollapragada et al. (2023) for augmented Lagrangian methods, and Bollapragada and Wild (2023) for quasi-Newton methods.
@@ A.2 Convergence of Stochastic Gradient Methods
We give a brief overview on the convergence results of SGD and adaptive gradient methods. The convergence of SGD for smooth nonconvex functions was first analyzed in Ghadimi and Lan (2013), assuming a uniform bound on the variance of stochastic gradients. Arjevani et al. (2023) gave a tight lower bound. Bottou et al. (2018) extended the convergence result with the so-called affine variance noise model. Different assumptions on the second moment of stochastic gradients appearing in the literature were reviewed in Khaled and Richtárik (2023), with a view to developing a general convergence theory of SGD for smooth nonconvex functions.
@@ Appendix B The Overall Algorithm
Sample an i.i.d. data batch $\mathcal{B}_{k}\subseteq\mathcal{D}$ uniformly from $\mathcal{D}$ of size $b_{k}$
@@
$\hskip 21.33955pt\mathsf{T}_{j}(x_{k};\mathcal{B}_{k},\eta)\coloneqq\frac{1}{b_{k}-1}\sum_{i\in\mathcal{B}_{k}}\left(\partial_{j}f(x_{k};\xi_{i})-\partial_{j}F_{\mathcal{B}_{k}}(x_{k})\right)^{2}\big{/}\left(\eta^{2}(\partial_{j}F_{\mathcal{B}_{k}}(x_{k}))^{2}\right),\;j\in\llbracket d\rrbracket$
@@ C.2 Preparatory Definitions, Propositions and Lemmas
This is part of the result of Bollapragada et al. (2018a), Lemma 3.1. We include its proof here with our notation for completeness.
@@ C.2.3 Technical Lemmas
We state without proof a similar descent lemma for $(L_{0},L_{1})$-smooth functions. Its proof can be found in Zhang et al. (2020a), Lemma A.3.
@@ C.3 Convergence Results for AdaGrad-Norm and AdaGrad
$\mathbb{E}_{k}[F(x_{k+1})]\leqslant F(x_{k})-\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)g_{k}\right]\right\rangle+\frac{L\alpha^{2}}{2}\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}.$ (C.5)
@@
$\frac{L\alpha^{2}}{2}\sum_{k=1}^{K}\mathbb{E}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}+\frac{\tau L^{2}\alpha^{3}}{2}(1+\rho\tau)\sum_{k=1}^{K}\mathbb{E}\left[\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}\right]\leqslant\frac{L\alpha^{2}}{2}(\mathbb{E}\log v_{K}-\log v_{0})+\tau L^{2}\alpha^{3}(1+\rho\tau)\frac{1}{\sqrt{v_{0}}}.$
@@
$\frac{\alpha}{2}\left(1-\frac{1}{\rho}\right)\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]\\ \leqslant F(x_{1})-\mathbb{E}[F(x_{K+1})]+\frac{\tau\alpha}{2}\mathbb{E}[\varphi_{0}-\varphi_{K}]+\frac{L\alpha^{2}}{2}(\mathbb{E}\log v_{K}-\log v_{0})+\tau L^{2}\alpha^{3}(1+\rho\tau)\frac{1}{\sqrt{v_{0}}},$
@@
By Lemma C.2, we have $\mathbb{E}[\sqrt{v_{k}}]\leqslant 2c_{1}\tau+8c_{2}\tau\log(c_{2}\tau+1)\eqqcolon c_{3}$. Since, by Hölder’s inequality,
$c_{1}+2c_{2}\log\mathbb{E}[\sqrt{v_{k}}]\geqslant\sum_{k=1}^{K}\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]\geqslant\mathbb{E}\left[\frac{1}{\sqrt{v_{K}}}\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}\right]\geqslant\frac{\mathbb{E}\left[\sqrt{\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}}\right]^{2}}{\mathbb{E}[\sqrt{v_{K}}]},$
@@
$\mathbb{E}\left[\sqrt{\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}}\right]^{\negthinspace 2}\leqslant\mathbb{E}[\sqrt{v_{K}}]\cdot\left(c_{1}+2c_{2}\log\mathbb{E}[\sqrt{v_{K}}]\right)\leqslant c_{3}(c_{1}+2c_{2}\log c_{3}).$
@@
As will be seen later, under the choice of the parameters $(\rho_{1},\rho_{2},\omega)$, the constant step size satisfies $\alpha\leqslant 1/L_{1}$ and $\left\lVert\nabla F(x_{k})/\sqrt{v_{k}}\right\rVert\leqslant 1$, we have
$\left\lVert x_{k+1}-x_{k}\right\rVert=\alpha\left\lVert\frac{\nabla F(x_{k})}{\sqrt{v_{k}}}\right\rVert\leqslant\frac{1}{L_{1}}.$
@@
$\quad\;-\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle+\frac{L_{1}\alpha^{2}\left\lVert\nabla F(x_{k})\right\rVert}{2}\cdot\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}$
@@
since $\alpha\leqslant 1/L_{1}$. By the A.M.-G.M. inequality (i.e., $2ab\leqslant\frac{1}{\rho_{1}}a^{2}+\rho_{1}b^{2}$ for any $\rho_{1}>0$), we have
@@
$\rho_{1}\alpha\tau\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right]\\ \leqslant\alpha\left(\frac{\rho_{1}}{\rho_{2}}+2\omega\right)\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{2L_{0}^{2}}{\sqrt{v_{k-1}}}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+\frac{\rho_{1}\rho_{2}L_{0}^{2}\tau^{2}}{\sqrt{v_{k-1}}}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+\rho_{1}\alpha\tau\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}].$